# On the Characteristics of the Conjugate Function Enabling Effective Dual Decomposition Methods

### v1

@@ Abstract
We investigate a novel characteristic of the conjugate function associated to a generic convex optimization problem which can subsequently be leveraged for efficient dual decomposition methods. In particular, under mild assumptions, we show a specific region of the domain of the conjugate function where there is always a ray originating from any point of the region such that the gradients of the conjugate remain constant along the ray. We refer to this characteristic as a fixed gradient over rays (FGOR). We further show that the dual function inherits from the conjugate the characteristic of FGOR. Then we provide a thorough exposition of applying the FGOR characteristic to dual subgradient methods. More importantly, we leverage FGOR to devise a simple stepsize rule that can be prepended with state-of-the-art stepsize methods enabling them to be more efficient. Furthermore, we investigate how the characteristic of FGOR is used when solving the global consensus problem. We show that FGOR can be exploited not only to expedite the convergence of the dual decomposition methods but also to reduce the communication overhead when distributed implementations are sought. Numerical experiments using quadratic objectives and a regularized linear regression with a real data set are conducted to compare the practical performance of FGOR with state-of-the-art stepsize methods. Results show that the proposed approach can significantly improve the convergence of existing methods while saving a considerable amount of communication  overhead.
Index Terms:
Distributed optimization, conjugate function, subgradient method, dual decomposition.
@@ I Introduction
Distributed algorithms for optimization problems have become necessary and pervasive in various application domains such as signal processing, machine learning, wireless communications, control systems, robotics, and many others [1, 2, 3, 4, 5, 6]. These problems are typically of very large scale since they deal with thousands, millions, or even more variables. In this respect, effective deployment of such algorithms requires an appeal to light communication among subsystems involved in solving the optimization problem and less computational effort per subsystem, and those inevitably raise questions about how fast the deployed algorithms converge. As such, in many real-world applications, the most commonly employed method to solve optimization problems is the (sub) gradient method. The choice of (sub) gradient method sparked numerous efforts by many researchers in designing various stepsize rules and applying them to yield faster convergences [7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35].
The existing literature that designs adaptive stepsizes primarily considers subgradient methods within the primal domain. These results can be readily applied to distributed algorithms that are based on duality [36, 37, 38, 39]. More specifically, in the case of dual decomposition algorithms for distributed optimization, the master problem that coordinates the subproblems at subsystems is often solved by using subgradient methods [40] where the earlier results for stepsizes can be adopted. However, these stepsize methods are applied to the dual domain without exploiting possible useful characteristics of such domain. Thus, we now pose the natural question: can we exploit any specific characteristics of the dual function itself to further scrutinize the design of stepsize rules that go beyond those designs within the primal domain for effective convergence of the subgradient method associated to the master problem?
@@
where the variable is $\mathbf{y}\in{\rm I\!R}^{n}$, $f_{0}:{\rm I\!R}^{n}\rightarrow{\rm I\!R}$ is the convex objective function, $\mathbf{A}\in{\rm I\!R}^{m\times n}$, $\mathbf{b}\in{\rm I\!R}^{m}$, and $\mathcal{Y}$ is a convex set in ${\rm I\!R}^{n}$. Equivalently, one may reformulate problem (1) as
@@
where $\delta_{\mathcal{Y}}$ denotes the indicator function of the set $\mathcal{Y}$, cf.  Definition 2. Note that in a decomposition setting, $f_{0}$ and constraints are endowed with structural properties such as separability, enabling the application of the dual decomposition [40].
The convergence of dual decomposition methods applied to problem (2) hinges predominantly on the characteristics of the problem’s associated dual function. Note that the dual function of (2) with $\mathbf{A}\mathbf{y}=\mathbf{b}$ being the constraint to which Lagrange multipliers are associated, is intimately connected with the conjugate function of $f$ [41, § 5.1.6]. As such, to answer our posed problem, we investigate a novel characteristic of the conjugate function of $f$. In particular, we show that the variational geometry of $\mathcal{Y}$, cf.  [42, Fig. 6-8] and nondifferentiable properties of $f_{0}$ induce a specific characteristic on the conjugate function. We effectively use this characteristic to devise a precursory yet simple stepsize rule that can be prepended with other stepsizes [7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35] while improving their convergences. The proposed precursory stepsize rule enables not only distributed but also efficient implementation of the algorithms with a light communication overhead.
@@ I-A Related Work
Employing dual decomposition techniques to solve a primal problem of the form (2) with adequate structural properties solely relies on solving the associated dual problem [40, § 2]. Recall that such a dual problem is often solved by using subgradient methods. The traditional stepsize methods [8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 7, 20, 21], that have already been designed for subgradient methods are constant stepsize, line search, polynomially decay stepsizes, and geometrically decay stepsizes. These stepsizes directly apply to solving the dual problem, provided the dual function conforms to necessary regularity conditions. We give a short overview of these schemes in the following.
@@
Given the dual function satisfies some gradient Lipschitz conditions and the Lipschitz constant is known, then a constant stepsize can be chosen appropriately to ensure convergence of the algorithms [7, § 1.4]. The concept of line search dates back to the 1960s [8, 9] and is used in numerous application domains [10, 11, 12]. It is a commonly used method for determining an appropriate stepsize, especially when the gradient Lipschitz constants are unknown. However, unlike the constant stepsize, employing line searches does not enable an effective distributed implementation of dual decomposition methods. For example, in the case of a backtracking line search, there will be an overwhelming communication overhead among subsystems even when implementing one iteration of the line search itself, cf.  [41, p. 464].
@@
The polynomially decay stepsizes such as $\gamma_{0}/k$ and $\gamma_{0}/\sqrt{k}$, where $\gamma_{0}>0$ is an appropriately chosen constant [7, § 5.3], [13, 14, 15] are commonplace and highly popular in distributed optimization. The reason is that they can be readily applied to respective distributed algorithms because of their simple dependency on the iteration index $k$. Nonetheless, their drawback lies in the fact that, as they progressively decrease with each iteration, they tend to yield slow convergence [43, 44].
@@
The geometrically decay stepsize strategy for subgradient methods originated from [16, 17, 18]. A comprehensive exposition is found in [19, § 2.3]. The resulting linear rate of convergences associated to such stepsizes has gained attention in recent studies [20, 21]. However, linear convergence is guaranteed under stringent conditions restricting the use of such stepsizes in practice [19, Theorem 2.7].
@@
The stepsize rules considered in first-order methods [22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35] are different from the direct application of any traditional stepsizes discussed above. More importantly, these stepsize strategies are designed to further improve the convergences, capture more general problem formulations, and handle problems when the gradient Lipschitz constants are unknown. The rules therein can be classified as step decay stepsize [22, 23, 24, 25, 26, 27] and running attributes-dependent stepsizes [35, 28, 29, 30, 31, 32, 33, 34]. We overview them in the following.
@@
The step decay stepsize is widely adopted in stochastic non-convex optimization [22, 23, 24] and is used as a potential model for training deep neural networks [25, 26]. This stepsize rule is characterized by maintaining constant stepsizes within stages of the algorithm iterations while decreasing it at each subsequent stage. The basic idea is to start stepping more aggressively in the initial stages and gradually become less aggressive in subsequent stages. The rule intrinsically includes two decisions: 1) how the stepsize is reduced and 2) how the stage length is changed, over the stage count. Typically, the rule reduces the stepsize by a fixed factor at each stage, i.e., the stepsize is $\alpha^{t}$, where $\alpha\in(0,1)$ and $t$ is the stage count. This policy, where the stepsize is reduced by a factor after a predetermined number of epochs, instead of at every iteration, is viewed as a variant of the geometrically decay stepsize. In contrast, the stepsize schedule employed in [24] is proportional to $1/t$, where $t$ is the stage count. On the other hand, stage lengths are changed either with linear growth [24] or exponential growth [22] or sometimes are kept constant [23, 27]. In general, almost all step decay stepsize rules can be employed for dual decomposition methods.
@@
The running attributes-dependent stepsizes that rely on algorithms’ runtime data, such as decision variables, gradients, and function values, are extensively studied in [28, 29, 30, 31, 32, 33, 34, 35]. Polyak’s stepsize [28] is an appealing rule demonstrating linear convergence rates for general strongly convex functions, even under non-differentiable settings. However, it requires the optimal value of the underlying problem or initial guess of it with subsequent refinement to compute the stepsize in every iteration. Despite such restrictions, recent studies have revived interest in Polyak’s stepsize due to its linear convergence properties, cf.  [29, 30, 31].
@@
Another running attributes-dependent stepsize is the Barzilai-Borwein (BB) rule [32, 33, 34]. Unlike Polyak’s stepsize, the BB rule doesn’t require the knowledge of the optimal value of the underlying problem. Initial BB results are limited to two-dimensional quadratic objective functions with linear convergence guarantees [32]. Linear convergence with BB stepsizes under strictly convex quadratic functions, in general, has been established in [33]. Extensions of similar convergence results under more general settings, in particular, for smooth and strongly convex functions have been established in [34], however under more restricted assumptions. In [35], a more appealing stepsize is proposed without the need for any restricted assumptions like those in Polyak’s stepsize and BB rule. More specifically, the stepsize proposed in [35] requires no specific assumptions beyond the local gradient Lipschitz continuity. In a dual decomposition setting, all the above stepsize rules  [28, 29, 30, 31, 32, 33, 34, 35] can be deployed, given the necessary runtime data in each iteration is communicated to the entity that is updating the dual variables, e.g., a parameter server.
@@
When solving the dual problem in a dual decomposition setting, one can always employ first-order algorithms [7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 24, 23, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35] that are devised for subgradient methods. Nevertheless, it is instructive to observe that the dual function of the underlying primal formulation, can be endowed with certain characteristics to design stepsize rules to further improve the convergences. In this respect, dualization of strong convexity [42, Prop. 12.60] which entangles the gradient Lipschitzian properties of primal functions and strong convexity of the associated dual functions and vice versa is one of the most classical results that enables the choice of constant stepsize length in a dual decomposition setting, cf.  [7, § 1.4]. But to the best of the authors’ knowledge, there seems to be no other explicit record previously of the employment of dual characteristics towards designing stepsize rules for dual decomposition methods.

### v3

@@ Abstract
We investigate a novel characteristic of the conjugate function associated to a generic convex optimization problem, which can subsequently be leveraged for efficient dual decomposition methods. In particular, under mild assumptions, we show that there is a specific region in the domain of the conjugate function such that for any point in the region, there is always a ray originating from that point along which the gradients of the conjugate remain constant. We refer to this characteristic as a fixed gradient over rays (FGOR). We further show that this characteristic is inherited by the corresponding dual function. Then we provide a thorough exposition of the application of the FGOR characteristic to dual subgradient methods. More importantly, we leverage FGOR to devise a simple stepsize rule that can be prepended with state-of-the-art stepsize methods enabling them to be more efficient. Furthermore, we investigate how the FGOR characteristic is used when solving the global consensus problem, a prevalent formulation in diverse application domains. We show that FGOR can be exploited not only to expedite the convergence of the dual decomposition methods but also to reduce the communication overhead. FGOR is extended to nonconvex formulations, and its advantages in stochastic optimization are demonstrated. Numerical experiments using quadratic objectives and a regularized least squares regression with real datasets are conducted. The results show that FGOR can significantly improve the performance of existing stepsize methods and outperform the state-of-the-art splitting methods on average in terms of both convergence behavior and communication efficiency.
@@ I Introduction
Distributed algorithms for optimization problems have become necessary and pervasive in various application domains such as signal processing, machine learning, wireless communications, control systems, robotics, and many others [Cao-2012-overview, di_2020, Nedic-Distributed-Optimization-for-Control-2018, Yang-survey-of-distributed-optimization-2019, Palomar-book-2010, Carlo-Machine-Learning, condat_2013_splitting_methods, Boyd-Parikh-Chu-Peleato-Eckstein-2010, He_2021, Bai_2024]. These problems are typically of very large scale since they deal with thousands, millions, or even more variables. In this respect, effective deployment of such algorithms requires an appeal to light communication among subsystems involved in solving the optimization problem and less computational effort per subsystem, and those inevitably raise questions about how fast the deployed algorithms converge. Although second-order algorithms such as Newton’s methods [Hintermuller_2010], [Boyd-Vandenberghe-04, § 10] exhibit fast convergence properties, they are generally less appealing for distributed large-scale optimization due to their substantial computational overhead (e.g., the need for large matrix inversions). As such, in many real-world applications, the most commonly employed method to solve optimization problems is the (sub) gradient method which is a first-order algorithm. The choice of (sub) gradient method sparked numerous efforts by many researchers in designing various stepsize rules and applying them to yield faster convergences [Polyak_Intro_Opt, goldstein_1962, armijo_1966, beck_2009, bello_2016, zhang2004nonmonotone, Asl_2020, khirirat_2023, Thinh_2018, Thinh_2021, zhor_1968_1, zhor_1968_2, goffin_1977, shor_2012, Damek_2018, zhu_2019, aybat_2019, wang_2021, chen_2019, krizhevsky_2012, He_2016, ge_2019, polyak_1969, hazan_2022, loizou_2021, wang_2023, barzilai_1988, dai_2002, burdakov_2019, Yura_2020].
The existing literature that designs adaptive stepsizes primarily considers subgradient methods within the primal domain. These results can be readily applied to distributed algorithms that are based on duality [Sindri_2020, R1-Liu-2021, Necoara-inexact-dual-decomposition-2014, Yifan-Asynchrony-and-Inexactness-2023]. More specifically, in the case of dual decomposition algorithms for distributed optimization, the master problem that coordinates the subproblems at subsystems is often solved by using subgradient methods [Boyd-EE364b-PrimDualDecomp-07] where the earlier results for stepsizes can be adopted. However, these stepsize methods are applied to the dual domain without exploiting possible useful characteristics of such domain. Thus, we now pose the natural question: can we exploit any specific characteristics of the dual function itself to further scrutinize the design of stepsize rules that go beyond those designs within the primal domain for effective convergence of the subgradient method associated to the master problem?
@@
where the variable is $\mathbf{y}\in{\rm I\!R}^{n}$, $f_{0}:{\rm I\!R}^{n}\rightarrow{\rm I\!R}$ is the convex objective function, $\mathcal{Y}$ is a convex set in ${\rm I\!R}^{n}$, and $\mathbf{b}\in{\rm I\!R}^{m}$. Without loss of generality, we assume $\mathbf{A}\in{\rm I\!R}^{m\times n}$ with $m<n$ and rank $\mathbf{A}=m$. Equivalently, one may reformulate problem (1) as
@@
where $\delta_{\mathcal{Y}}$ denotes the indicator function of the set $\mathcal{Y}$, cf.  Definition 2. Note that in a decomposition setting, $f_{0}$ and constraints are endowed with structural properties such as separability, enabling the application of the dual decomposition [Boyd-EE364b-PrimDualDecomp-07].
The convergence of dual decomposition methods applied to problem (2) hinges predominantly on the characteristics of the problem’s associated dual function. Note that the dual function of (2) with $\mathbf{A}\mathbf{y}=\mathbf{b}$ being the constraint to which Lagrange multipliers are associated, is intimately connected with the conjugate function of $f$ [Boyd-Vandenberghe-04, § 5.1.6]. As such, to answer our posed problem, we investigate a novel characteristic of the conjugate function of $f$. In particular, we show that the variational geometry of $\mathcal{Y}$, cf.  [Rockafellar-98, Fig. 6-8] and nondifferentiable properties of $f_{0}$ induce a specific characteristic on the conjugate function. We effectively use this characteristic to devise a precursory yet simple stepsize rule that can be prepended with other stepsizes [Polyak_Intro_Opt, goldstein_1962, armijo_1966, beck_2009, bello_2016, zhang2004nonmonotone, Asl_2020, khirirat_2023, Thinh_2018, Thinh_2021, zhor_1968_1, zhor_1968_2, goffin_1977, shor_2012, Damek_2018, zhu_2019, aybat_2019, wang_2021, chen_2019, krizhevsky_2012, He_2016, ge_2019, polyak_1969, hazan_2022, loizou_2021, wang_2023, barzilai_1988, dai_2002, burdakov_2019, Yura_2020] while improving their convergences. The proposed precursory stepsize rule enables not only distributed but also efficient implementation of the algorithms with a light communication overhead. Our exposition begins with convex problems, as they provide a fundamental basis for developing insights that enable extensions to more general settings (e.g., nonconvex problems and stochastic methods). Furthermore, in the machine learning domain, these convex formulations correspond to simple and computationally lightweight models, such as linear/logistic regression and support vector machines, and are also appealing from an explainability and interpretability standpoint, properties that are particularly important in sensitive application domains such as medicine and bioinformatics  [karim2023explainable, karim2022interpreting].
@@ I-A Related Work
Employing dual decomposition techniques to solve a primal problem of the form (2) with adequate structural properties solely relies on solving the associated dual problem [Boyd-EE364b-PrimDualDecomp-07, § 2]. Recall that such a dual problem is often solved by using subgradient methods. The traditional stepsize methods [goldstein_1962, armijo_1966, beck_2009, bello_2016, Asl_2020, khirirat_2023, Thinh_2018, Thinh_2021, zhor_1968_1, zhor_1968_2, goffin_1977, shor_2012, Polyak_Intro_Opt, Damek_2018, zhu_2019], that have already been designed for subgradient methods are constant stepsize, line search, polynomially decay stepsizes, and geometrically decay stepsizes. These stepsizes directly apply to solving the dual problem, provided the dual function conforms to necessary regularity conditions. We give a short overview of these schemes in the following.
@@
Given the dual function satisfies some gradient Lipschitz conditions and the Lipschitz constant is known, then a constant stepsize can be chosen appropriately to ensure convergence of the algorithms [Polyak_Intro_Opt, § 1.4]. The concept of line search dates back to the 1960s [goldstein_1962, armijo_1966] and is used in numerous application domains [beck_2009, bello_2016, zhang2004nonmonotone, Asl_2020]. It is a commonly used method for determining an appropriate stepsize, especially when the gradient Lipschitz constants are unknown. However, unlike constant stepsize approaches, incorporating line search techniques does not lend itself to an efficient distributed implementation of dual decomposition methods. This limitation arises because, in each iteration of the dual decomposition algorithm, additional Lagrangian minimization steps must be coordinated among subsystems to compute the corresponding dual function and gradient values, which are subsequently used to verify the line search conditions. Furthermore, implementing even a single iteration of the line search procedure introduces significant signaling overhead among subsystems [Boyd-Vandenberghe-04, p. 464][zhang2004nonmonotone, p. 1044].
@@
The polynomially decay stepsizes such as $\gamma_{0}/k$ and $\gamma_{0}/\sqrt{k}$, where $\gamma_{0}>0$ is an appropriately chosen constant [Polyak_Intro_Opt, § 5.3], [khirirat_2023, Thinh_2018, Thinh_2021] are commonplace and highly popular in distributed optimization. The reason is that they can be readily applied to respective distributed algorithms because of their simple dependency on the iteration index $k$. Nonetheless, their drawback lies in the fact that, as they progressively decrease with each iteration, they tend to yield slow convergence [Bazaraa_1981, Bianchi_2022].
@@
The geometrically decay stepsize strategy for subgradient methods originated from [zhor_1968_1, zhor_1968_2, goffin_1977]. A comprehensive exposition is found in [shor_2012, § 2.3]. The resulting linear rate of convergences associated to such stepsizes has gained attention in recent studies [Damek_2018, zhu_2019]. However, linear convergence is guaranteed under stringent conditions restricting the use of such stepsizes in practice [shor_2012, Theorem 2.7].
@@
The stepsize rules considered in first-order methods [aybat_2019, wang_2021, chen_2019, krizhevsky_2012, He_2016, ge_2019, polyak_1969, hazan_2022, loizou_2021, wang_2023, barzilai_1988, dai_2002, burdakov_2019, Yura_2020] are different from the direct application of any traditional stepsizes discussed above. More importantly, these stepsize strategies are designed to further improve the convergences, capture more general problem formulations, and handle problems when the gradient Lipschitz constants are unknown. The rules therein can be classified as step decay stepsize [aybat_2019, wang_2021, chen_2019, krizhevsky_2012, He_2016, ge_2019] and running attributes-dependent stepsizes [Yura_2020, polyak_1969, hazan_2022, loizou_2021, wang_2023, barzilai_1988, dai_2002, burdakov_2019]. We overview them in the following.
@@
The step decay stepsize is widely adopted in stochastic non-convex optimization [aybat_2019, wang_2021, chen_2019] and is used as a potential model for training deep neural networks [krizhevsky_2012, He_2016]. This stepsize rule is characterized by maintaining constant stepsizes within stages of the algorithm iterations while decreasing it at each subsequent stage. The basic idea is to start stepping more aggressively in the initial stages and gradually become less aggressive in subsequent stages. The rule intrinsically includes two decisions: 1) how the stepsize is reduced and 2) how the stage length is changed, over the stage count. Typically, the rule reduces the stepsize by a fixed factor at each stage, i.e., the stepsize is $\alpha^{t}$, where $\alpha\in(0,1)$ and $t$ is the stage count. This policy, where the stepsize is reduced by a factor after a predetermined number of epochs, instead of at every iteration, is viewed as a variant of the geometrically decay stepsize. In contrast, the stepsize schedule employed in [chen_2019] is proportional to $1/t$, where $t$ is the stage count. On the other hand, stage lengths are changed either with linear growth [chen_2019] or exponential growth [aybat_2019] or sometimes are kept constant [wang_2021, ge_2019]. In general, almost all step decay stepsize rules can be employed for dual decomposition methods.
@@
The running attributes-dependent stepsizes that rely on algorithms’ runtime data, such as decision variables, gradients, and function values, are extensively studied in [polyak_1969, hazan_2022, loizou_2021, wang_2023, barzilai_1988, dai_2002, burdakov_2019, Yura_2020]. Polyak’s stepsize [polyak_1969] is an appealing rule demonstrating linear convergence rates for general strongly convex functions, even under non-differentiable settings. However, it requires the optimal value of the underlying problem or initial guess of it with subsequent refinement to compute the stepsize in every iteration. Despite such restrictions, recent studies have revived interest in Polyak’s stepsize due to its linear convergence properties, cf.  [hazan_2022, loizou_2021, wang_2023].
@@
Another running attributes-dependent stepsize is the Barzilai-Borwein (BB) rule [barzilai_1988, dai_2002, burdakov_2019]. Unlike Polyak’s stepsize, the BB rule doesn’t require the knowledge of the optimal value of the underlying problem. Initial BB results are limited to two-dimensional quadratic objective functions with linear convergence guarantees [barzilai_1988]. Linear convergence with BB stepsizes under strictly convex quadratic functions, in general, has been established in [dai_2002]. Extensions of similar convergence results under more general settings, in particular, for smooth and strongly convex functions have been established in [burdakov_2019], however under more restricted assumptions. In [Yura_2020], a more appealing stepsize is proposed without the need for any restricted assumptions like those in Polyak’s stepsize and BB rule. More specifically, the stepsize proposed in [Yura_2020] requires no specific assumptions beyond the local gradient Lipschitz continuity. In a dual decomposition setting, all the above stepsize rules  [polyak_1969, hazan_2022, loizou_2021, wang_2023, barzilai_1988, dai_2002, burdakov_2019, Yura_2020] can be deployed, given the necessary runtime data in each iteration is communicated to the entity that is updating the dual variables, e.g., a parameter server.
@@
When solving the dual problem in a dual decomposition setting, one can always employ first-order algorithms [Polyak_Intro_Opt, goldstein_1962, armijo_1966, beck_2009, bello_2016, Asl_2020, khirirat_2023, Thinh_2018, Thinh_2021, zhor_1968_1, zhor_1968_2, goffin_1977, shor_2012, Damek_2018, zhu_2019, aybat_2019, chen_2019, wang_2021, krizhevsky_2012, He_2016, ge_2019, polyak_1969, hazan_2022, loizou_2021, wang_2023, barzilai_1988, dai_2002, burdakov_2019, Yura_2020] that are devised for subgradient methods. Nevertheless, it is instructive to observe that the dual function of the underlying primal formulation, can be endowed with certain characteristics to design stepsize rules to further improve the convergences. In this respect, dualization of strong convexity [Rockafellar-98, Prop. 12.60] which entangles the gradient Lipschitzian properties of primal functions and strong convexity of the associated dual functions and vice versa is one of the most classical results that enables the choice of constant stepsize length in a dual decomposition setting, cf.  [Polyak_Intro_Opt, § 1.4]. But to the best of the authors’ knowledge, there seems to be no other explicit record previously of the employment of dual characteristics towards designing stepsize rules for dual decomposition methods.